The least of three consecutive integers is divided by 10, the next is divided by 17,the greatest is divided by 26. What are the numbers if the sum of the quotients is 10?
Answer by ewatrrr(24785)


Hi
Let x,(x+1),(x+2) represent the 3 consecutive numbers
Question states***
x/10 + (x+1)/17 +(x+2)/26 = 10
Solving for x
17*26x + 10*26(x+1)+ 10*17(x+2)= 10*10*17*26
442x + 260x + 170x + 260 + 340 = 44200
872x = 43600
x = 50 The 3 consecutive numbers are 50,51,52
CHECKING our Answer***
5 + 3 + 2 = 10